# Hydrostatic Newton-Cartan Membranes

### v1

@@ V Young-Laplace equation for axisymmetric solutions
The Young-Laplace equation takes the form
$2\left(a^{(2)}_{1}+a_{2}^{(2)}\right)D^{2}K_{1}+\left(\beta^{(1)}_{1}+2\beta^{(2)}_{1}K_{1}\right)\bar{a}^{2}-\left(n^{(1)}_{1}+2n^{(2)}_{1}K_{1}\right)\left(D\cdot a\right)$ (138)
$+2\beta^{(2)}_{2}\left(K_{cd}a^{c}a^{d}\right)-2n^{(2)}_{2}\left(K_{cd}h^{ac}D_{a}a^{d}\right)-4\left(n^{(2)}_{1}+n^{(2)}_{2}\right)a^{c}D_{c}K_{1}$
$-\chi K_{1}-a_{1}^{(1)}K_{1}^{2}+a_{1}^{(1)}K_{2}-\left(a_{1}^{(2)}+a_{2}^{(2)}\right)\left(K_{1}^{2}-2K_{2}\right)K_{1}$
$+nk_{0}+n_{1}^{(1)}K_{1}k_{0}+n_{1}^{(2)}K^{2}_{1}k_{0}+n_{2}^{(2)}K_{2}k_{0}$ $=0.$
From the form of the Young-Laplace equation we can notice that it is convenient to change the transport coefficients from the set $\{\chi,a^{(1)}_{1},a^{(2)}_{1},a^{(2)}_{2}\}$ to the set $\{\chi,a^{(1)}_{1},a^{(2)}_{2},\kappa\}$, where $\kappa$ is defined through
$a^{(2)}_{1}$ $=\kappa-a_{2}^{(2)}.$ (139)
The corresponding derivatives of $\kappa$ with respect to $\mu$ are related to $n^{(2)}_{1},n^{(2)}_{2},\beta^{(2)}_{1}$ and $\beta^{(2)}_{2}$ through
$n^{(2)}_{1}$ $=n_{\kappa}-n_{2}^{(2)}\,,$ (140)
$\beta^{(2)}_{1}$ $=\beta_{\kappa}-\beta_{2}^{(2)}.$ (141)
Under this redefinition, the Young-Laplace equation becomes the following
$2\kappa D^{2}K_{1}+\left(\beta^{(1)}_{1}+2\left(\beta_{\kappa}-\beta^{(2)}_{2}\right)K_{1}\right)\bar{a}^{2}-\left(n^{(1)}_{1}+2\left(n_{\kappa}-n_{2}^{(2)}\right)K_{1}\right)\left(D\cdot a\right)$ (142)
$+2\beta^{(2)}_{2}\left(K_{cd}a^{c}a^{d}\right)-2n^{(2)}_{2}\left(K_{cd}h^{ac}D_{a}a^{d}\right)-4n_{\kappa}a^{c}D_{c}K_{1}$
$-\chi K_{1}a_{1}^{(1)}\left(K_{2}-K_{1}^{2}\right)-\kappa\left(K_{1}^{2}-2K_{2}\right)K_{1}$
$+nk_{0}+n_{1}^{(1)}K_{1}k_{0}+n_{2}^{(2)}\left(K_{2}-K^{2}_{1}\right)k_{0}+n_{\kappa}K_{1}^{2}k_{0}$ $=0.$
For simplicity we will restate the Young-Laplace equation as
$Y+A_{\beta}+N^{(1)}_{1}+N^{(2)}_{2}+N_{\kappa}-\frac{n}{4}\omega^{2}\rho\sin\psi=0,$ (143)
where $Y$ contain the terms that do not involve the vorticity explicitly
$Y$ $\equiv 2\kappa D^{2}K_{1}-\chi K_{1}+a_{1}^{(1)}\left(K_{2}-K^{2}_{1}\right)-\kappa\left(K^{2}_{1}-2K_{2}\right)K_{1}$ (144)
$=-\frac{2a_{1}^{(1)}\psi^{\prime}\cos\psi\sin\psi}{\rho}+\chi\left(\psi^{\prime}\cos\psi+\frac{\sin\psi}{\rho}\right)$
$+\kappa\left[2\psi^{\prime\prime\prime}\cos^{2}\psi-8\psi^{\prime\prime}\psi^{\prime}\cos^{2}\psi\sin\psi+\frac{4\psi^{\prime\prime}\cos^{3}\psi}{\rho}+2\left(\psi^{\prime}\right)^{3}\cos\psi\sin^{2}\psi\right]$
$+\kappa\left[-\cos^{3}\psi\left(\psi^{\prime}\right)^{3}-\frac{7\left(\psi^{\prime}\right)^{2}\cos^{2}\psi\sin\psi}{\rho}+\frac{\psi^{\prime}\sin^{2}\psi\cos\psi}{\rho^{2}}+\frac{2\cos^{2}\psi\sin\psi}{\rho^{3}}\right]$
$+\kappa\left[\frac{\sin^{3}\psi}{\rho^{3}}-\frac{2\psi^{\prime}\cos^{3}\psi}{\rho^{2}}+\frac{\psi^{\prime}\sin^{2}\psi\cos\psi}{\rho^{2}}\right],$
$A_{\beta}$ contains the terms related to $\beta^{(1)}_{1},\beta^{(1)}_{2},$ and $\beta_{\kappa}$
$A_{\beta}$ $=\beta_{1}^{(1)}\left(\frac{\omega^{4}}{16}\rho^{2}\cos^{2}\psi\right)+\beta_{2}^{(2)}\left(\frac{\omega^{4}}{8}\rho\cos^{2}\psi\sin\psi\right)$ (145)
$-\beta_{\kappa}\left(\frac{\omega^{4}}{8}\rho^{2}\cos^{2}\psi\right)\left(\psi^{\prime}\cos\psi+\frac{\sin\psi}{\rho}\right),$
$N^{(I)}_{i}$ contains terms associate to $n^{(I)}_{i}$
$N^{(1)}_{1}$ $=\frac{\omega^{2}}{8}n^{(1)}_{1}\left(3+\cos\left(2\psi\right)\right)\,,$ (146)
$N^{(2)}_{2}$ $=\frac{\omega^{2}}{2}n_{2}^{(2)}\cos^{2}\psi\left(\psi^{\prime}\cos\psi+\frac{\sin\psi}{\rho}\right),$ (147)
and $N_{\kappa}$ contains terms associated to $n_{\kappa}$
$N_{\kappa}$ $=-\frac{\omega^{2}}{4\rho}n_{\kappa}\sin^{3}\psi-2\omega^{2}n_{\kappa}\psi^{\prime}\cos^{3}\psi+\frac{5}{4}\omega^{2}n_{\kappa}\rho\left(\psi^{\prime}\right)^{2}\cos^{2}\psi\sin\psi-\omega^{2}n_{\kappa}\rho\psi^{\prime\prime}\cos^{3}\psi.$ (148)
[…]

### v2

@@ 5 Young-Laplace equation for axisymmetric solutions for Canham-Helfrich-like membranes
Combining (134) and (136), we find that the Young-Laplace equation takes the form
$2\left(a^{(2)}_{1}+a_{2}^{(2)}\right)D^{2}K+\left(\beta^{(1)}_{1}+2\beta^{(2)}_{1}K\right)\bar{a}^{2}-\left(n^{(1)}_{1}+2n^{(2)}_{1}K\right)\left(D\cdot a\right)$ (138)
$+2\beta^{(2)}_{2}\left(K_{cd}a^{c}a^{d}\right)-2n^{(2)}_{2}\left(K_{cd}h^{ac}D_{a}a^{d}\right)-4\left(n^{(2)}_{1}+n^{(2)}_{2}\right)a^{c}D_{c}K$
$-\chi K-a_{1}^{(1)}K^{2}+a_{1}^{(1)}\left(K\cdot K\right)-\left(a_{1}^{(2)}+a_{2}^{(2)}\right)\left(K^{2}-2\left(K\cdot K\right)\right)K$
$-n\left(a\cdot n\right)-n_{1}^{(1)}K\left(a\cdot n\right)-n_{1}^{(2)}K^{2}\left(a\cdot n\right)-n_{2}^{(2)}\left(K\cdot K\right)\left(a\cdot n\right)$ $=0.$
To understand the possible applications and interpretations of the solutions of (138), we can recall that, in the context of biophysics, one possibility is to interpret the partition function (129) as the energy functional of a biophysical membrane [3]. In particular, we can relate the coefficients $\{\chi,a^{(1)}_{1},a^{(2)}_{1},a^{(2)}_{2}\}$ to those appearing on the Canham–Helfrich free energy model for lipid vesicles. The Canham–Helfrich free energy has the typical form [23]
$\mathcal{F}_{CH}=\int d^{3}x\left[\gamma+\frac{\kappa}{2}\left(K-K_{0}\right)^{2}+\bar{\kappa}\mathcal{R}\right],$ (139)
where $\mathcal{R}=h^{ac}\mathcal{R}^{d}{}_{adc}$ is the spatial Ricci scalar. The coefficients that appear in (139) are the surface tension $\gamma$, the bending modulus $\kappa$, the spontaneous curvature $K_{0}$, and the Gaussian modulus $\bar{\kappa}$. The partition function (129) can be related to the free energy functional (139) through the identifications
$\chi$ $=\gamma+\frac{\kappa K_{0}^{2}}{2},$ (140)
$a^{(1)}_{1}$ $=-\kappa K_{0},$ (141)
$a^{(2)}_{1}$ $=\frac{\kappa}{2}+\bar{\kappa},$ (142)
$a^{(2)}_{2}$ $=-\bar{\kappa},$ (143)
where we used the identity $\mathcal{R}=K^{2}-(K\cdot K)$. Using the Canham-Helfrich coefficients, the Young-Laplace equation (138) takes the form
$-K\gamma+\kappa\left[D^{2}K-\frac{1}{2}\left(K-K_{0}\right)^{2}K+\left(K-K_{0}\right)\left(K\cdot K\right)\right]-n_{\gamma}\left(a\cdot n\right)$
$+\kappa n_{0}\left[\left(D\cdot a\right)+\left(a\cdot n\right)\left(K-K_{0}\right)\right]-2n_{0}n_{\kappa}\bar{a}^{2}$
$-n_{\kappa}\left[2a^{m}D_{m}K+\frac{1}{2}\left(K-K_{0}\right)^{2}\left(a\cdot n\right)+\left(D\cdot a\right)\left(K-K_{0}\right)\right]-\beta_{0}\kappa\bar{a}^{2}$
$+\beta_{\kappa}\left(K-K_{0}\right)\bar{a}^{2}+n_{\bar{\kappa}}\left[2K_{cr}h^{rs}D_{s}a^{c}-2K\left(D\cdot a\right)+\left(a\cdot n\right)\left((K\cdot K)-K^{2}\right)\right]$
$+2\beta_{\bar{\kappa}}\left(K\bar{a}^{2}-K_{cd}a^{c}a^{d}\right)=\Delta p,$ (144)
where the coefficients $\{n_{\gamma},n_{0},n_{\kappa},n_{\bar{\kappa}},\beta_{0},\beta_{\kappa},\beta_{\bar{\kappa}}\}$ are defined as
$n_{0}$ $=\frac{\partial K_{0}}{\partial\mu},$ $n_{\kappa}$ $=\frac{\partial\kappa}{\partial\mu},$ $n_{\bar{\kappa}}$ $=\frac{\partial\bar{\kappa}}{\partial\mu},$ $n_{\gamma}=\frac{\partial\gamma}{\partial\mu},$
$\beta_{0}$ $=\frac{\partial^{2}K_{0}}{\partial\mu^{2}}\,$, $\beta_{\kappa}$ $=\frac{\partial^{2}\kappa}{\partial\mu^{2}},$ $\beta_{\bar{\kappa}}$ $=\frac{\partial^{2}\bar{\kappa}}{\partial\mu^{2}}.$ (145)
Notice that in (5) we included the possibility of having a pressure difference $\Delta p$ between the fluid outside and inside the membrane. It is known that the pressure difference will enter the Young-Laplace equation as shown in (5), see [3] for a derivation of this particular contribution using the Newton-Cartan sub-manifolds formalism. For completeness, we will now use the axisymmetric configuration discussed in section IV.B to write the explicit form of the differential Young-Laplace equation (5) in terms of the function $\phi(\rho)=\sin\psi(\rho)$
$\kappa\left(\phi^{2}-1\right)\phi^{\prime\prime\prime}+\kappa\left(\phi\phi^{\prime}+\frac{2(\phi^{2}-1)}{\rho}\right)\phi^{\prime\prime}-\frac{\kappa}{2}\left(\phi^{\prime 2}-\frac{3\phi\phi^{\prime}}{\rho}+\frac{3\phi^{2}-2}{\rho^{2}}\right)\phi^{\prime}$
$+\frac{\kappa}{2}\left(\frac{\phi^{2}-2}{\rho^{3}}\right)\phi+\gamma\left(\phi^{\prime}+\frac{\phi}{\rho}\right)+\frac{K_{0}\kappa}{2}\left(K_{0}+\frac{4\phi}{\rho}\right)\phi^{\prime}+\frac{\kappa K_{0}^{2}}{2}\frac{\phi}{\rho}$
$+\frac{n_{0}\kappa\omega^{2}}{4}\left(\rho\phi(K_{0}+2\phi^{\prime})+3\phi^{2}-2\right)+\frac{n_{\bar{\kappa}}\omega^{2}}{2}\left[(3\phi^{2}-1)\phi^{\prime}+\frac{(\phi^{2}-1)\phi}{\rho}\right]$
$+\frac{n_{\gamma}\omega^{2}}{4}\rho\phi+n_{\kappa}\omega^{2}\left[\frac{\rho(\phi^{2}-1)\phi^{\prime\prime}}{2}+\frac{\phi^{\prime}}{8}\left(\rho\phi(3\phi^{\prime}+4K_{0})+12\phi^{2}-8\right)\right$.
$\left.+\frac{\phi}{8}\left(\frac{\phi^{2}}{\rho}+6K_{0}\phi+\rho K_{0}^{2}\right)-\frac{K_{0}}{2}\right]+\frac{\omega^{4}n_{0}n_{\kappa}}{8}\rho^{2}(\phi^{2}-1)$
$+\frac{\omega^{4}\beta_{\kappa}\rho}{16}\left(\rho\phi^{2}(K_{0}+\phi^{\prime})-\rho(K_{0}+\phi^{\prime})+\phi^{3}-\phi\right)+\frac{\omega^{4}\beta_{\bar{\kappa}}}{8}\rho\phi(\phi^{2}-1)$
$+\frac{\omega^{4}\beta_{0}\kappa\rho^{2}(\phi^{2}-1)}{16}=\Delta p.$ (146)
Equation (5), for $\omega=0$ and constant thermodynamic coefficients $\{\gamma,\kappa,K_{0},\bar{\kappa}\}$, has been known for quite some time [23]. Under those conditions, equation (5) admits some analytic solutions such as the sphere, the torus, and the biconcave disc [29]. Note that when $\bar{\kappa}$ is taken to be constant, it does not contribute to the Young-Laplace equation because its associated term in the equilibrium partition function becomes a constant term determined by the topology of the membrane [^11] Gauss-Bonnet theorem tells us that in two spatial dimensions, the curvature $\mathcal{R}$ is a total derivative, whose integral over the spatial dimensions is the Euler characteristic, which is fully determined by the topology of the surface. If the coefficient in front of the $\mathcal{R}$ term in the equilibrium partition function is a constant, then its contribution can be integrated out, and it will therefore not contribute to the Young-Laplace equation.. In the rest of the section, we will look at some solutions to (5) assuming different functional dependencies for the thermodynamic coefficients to explore some of the shapes the equation allows. In what follows, we are going to need the equations of state for the thermodynamic coefficients. We will assume all of them are regular functions of the chemical potential $\mu$ and admit the generic expansion
$\gamma(\mu)$ $=\gamma^{(0)}+n_{\gamma}^{(0)}\mu+\frac{1}{2}\beta_{\gamma}^{(0)}\mu^{2}+$..., (147)
$\kappa(\mu)$ $=\kappa^{(0)}+n_{\kappa}^{(0)}\mu+\frac{1}{2}\beta_{\kappa}^{(0)}\mu^{2}+$..., (148)
$K_{0}(\mu)$ $=K_{0}^{(0)}+n_{0}^{(0)}\mu+\frac{1}{2}\beta_{0}^{(0)}\mu^{2}+$..., (149)
$\bar{\kappa}(\mu)$ $=\bar{\kappa}^{(0)}+n_{\bar{\kappa}}^{(0)}\mu+\frac{1}{2}\beta_{\bar{\kappa}}^{(0)}\mu^{2}+$..., (150)
where $\{\gamma^{(0)},n^{(0)}_{\gamma},\beta_{\gamma}^{(0)},\kappa^{(0)},n_{\kappa}^{(0)},\beta_{\kappa}^{(0)},K_{0}^{(0)},n_{0}^{(0)},\beta_{0}^{(0)},\bar{\kappa}^{(0)},$ $n_{\bar{\kappa}}^{(0)},\beta_{\bar{\kappa}}^{(0)}\}$ are all constants  [^12] In principle, these coefficients can be functions of the temperature. However, in equilibrium, the temperature must necessarily be constant and where “…” denotes higher order terms in the chemical potential that will eventually be taken to be zero and will not play any role in our discussion.
We will be interested in looking at closed membrane solutions. For solutions with the topology of the sphere, we will have the boundary conditions $\phi(0)=0$ and $\phi(r)=-1$, where $r$ is the maximum width of the membrane as measured in the ambient space and where the condition at $\rho=0$ ensures the solution has no kinks. For toroidal solutions, the boundary conditions will be $\phi(r_{+}-r_{-})=1$ and $\phi(r_{+}+r_{-})=-1$, where $r_{-}$ is the torus minor radius and $r_{+}$ is the torus major radius.
For all solutions that we looked at in this work, we have assumed thermodynamic relations linear in $\mu$, namely only $\{\gamma^{(0)},n^{(0)}_{\gamma},\kappa^{(0)},n^{(0)}_{\kappa},K^{(0)}_{0},n^{(0)}_{0}$, $\bar{\kappa}^{(0)},n^{(0)}_{\bar{\kappa}}\}$ have the possibility of being non-vanishing.
5.1 Spherical Solution
We will denote a spherical solution as $\phi_{S}(\rho)$. Such solution is given by
$\phi_{S}(\rho)=-\frac{\rho}{r},$ (151)
with $r$ the radius of the sphere. The function (151) is a solution to (5) when
$\gamma(\mu)$ $=\gamma^{(0)}+\left[-\frac{K^{(0)}_{0}n^{(0)}_{\kappa}}{2}+\frac{3K^{(0)}_{0}n^{(0)}_{\kappa}}{r}\right$.
[…]
$\kappa(\mu)$ $=\kappa^{(0)}+n^{(0)}_{\kappa}\mu,$ (153)
$K_{0}(\mu)$ $=K^{(0)}_{0}$, (154)
$\bar{\kappa}(\mu)$ $=\bar{\kappa}^{(0)}+n^{(0)}_{\bar{\kappa}}\mu,$ (155)
and when the following constraint equation is satisfied
$r^{2}\Delta\tilde{p}+r\left(\left(K^{(0)}_{0}\right)^{2}\kappa^{(0)}+2\tilde{\gamma}^{(0)}\right)-2K^{(0)}_{0}\kappa^{(0)}=0,$ (156)
where we defined the shifted pressure $\Delta\tilde{p}$ and the shifted surface tension $\tilde{\gamma}^{(0)}$ as
$\Delta\tilde{p}$ $=\Delta p+\frac{1}{2}K^{(0)}_{0}n_{\kappa}^{(0)}\omega^{2},$ (157)
$\tilde{\gamma}^{(0)}$ $=\gamma^{(0)}-\frac{1}{2}\left(n^{(0)}_{\kappa}+n_{\bar{\kappa}}^{(0)}\right)\omega^{2}.$ (158)
Notice that the sphere is a solution for $\omega=0$ [3]. For a rotating membrane, the Young-Laplace equation (5) admits a spherical solution under specific conditions. Given a linear dependence on $\mu$ of the elastic moduli, and a constant intrinsic curvature $K_{0}^{(0)}$, the surface tension relation (152) will determine the sphere radius, as long as the pressure difference is set according to (156).
5.2 Biconcave Disc Solution
We will denote a biconcave disc solution as $\phi_{D}(\rho)$. Such solution is given by
$\phi_{D}(\rho)=-\rho\left(\frac{1-rK^{(0)}_{0}\ln r}{r}+K^{(0)}_{0}\ln\rho\right),$ (159)
where $r$ is the maximum horizontal length attained by the discoid. The function (159) is a solution to (5) when
$\gamma(\mu)$ $=0,$ (160)
$\kappa(\mu)$ $=\kappa^{(0)}+n_{\kappa}^{(0)}\mu,$ (161)
$K_{0}(\mu)$ $=K^{(0)}_{0}$, (162)
$\bar{\kappa}(\mu)$ $=\bar{\kappa}^{(0)}-n_{\kappa}^{(0)}\mu,$ (163)
and where the pressure difference takes the value
$\Delta p=\frac{n_{\kappa}^{(0)}\omega^{2}}{2}.$ (164)
As in the case of $\omega=0$, the discoid is a solution of zero surface tension [3], but in contrast, it needs a nonzero pressure difference proportional to $\omega^{2}$. What seems very restrictive in the discoid solution is the relations for the elastic moduli, where their linear rate of change with $\mu$ differs only in their sign.
5.3 Torus Solution
We will denote a toroidal solution as $\phi_{T}(\rho)$
$\phi_{T}(\rho)=-\frac{\rho}{r_{-}}+\frac{r_{+}}{r_{-}}\,$, (165)
where $r_{-}$ is the minor radius of the torus and $r_{+}$ is the mayor radius of the torus. The function (165) is a solution to (5) when
$\gamma(\mu)$ $=\left[\frac{K^{(0)}_{0}\kappa^{(0)}\left(4-K^{(0)}_{0}r_{-}\kappa^{(0)}\right)}{2r_{-}}\right]$ (166)
$-\left[\frac{n^{(0)}_{\bar{\kappa}}\left(2-3K^{(0)}_{0}r_{-}\right)\omega^{2}}{6\left(K^{(0)}_{0}r_{-}-2\right)}\right]$
$-\frac{4n_{\bar{\kappa}}^{(0)}\left(K^{(0)}_{0}r_{-}-1\right)}{3r_{-}^{2}}\mu\,$,
$\kappa(\mu)$ $=\kappa^{(0)}+\frac{8n_{\bar{\kappa}}^{(0)}}{3\left(K^{(0)}_{0}r_{-}-2\right)}\mu\,$, (167)
$K_{0}(\mu)$ $=K^{(0)}_{0}\,$, (168)
$\bar{\kappa}(\mu)$ $=\bar{\kappa}^{(0)}+n^{(0)}_{\bar{\kappa}}\mu$ (169)
where $r_{-}$ and $r_{+}$ are not independent and are related through
$r_{+}=\sqrt{2}r_{-}\,$, (170)
and where the pressure is constrained to take the value
$\Delta p=\frac{2}{3r_{-}^{3}}\left[\frac{n^{(0)}_{\bar{\kappa}}r_{-}\left(3-2K^{(0)}_{0}r_{-}\right)\omega^{2}}{K^{(0)}_{0}r_{-}-2}-3K^{(0)}_{0}\kappa^{(0)}\right]\,.$ (171)
The torus is the third solution that is also a solution at $\omega=0$, with its radii that relate through (170). Nonetheless, the dependence of both $\gamma$ and $\kappa$ on $r_{-}$, imposes severe restrictions on the membrane material properties.
5.4 Biconcave disc compatible membranes
We will now look at some particular numerical solutions to (5). Solutions with the topology of the sphere and with thermodynamic coefficients with at most a linear dependence on the chemical potential, namely
$\gamma(\mu)$ $=\gamma^{(0)}+n^{(0)}_{\gamma}\mu\,$, (172)
$\kappa(\mu)$ $=\kappa^{(0)}+n^{(0)}_{\kappa}\mu\,$, (173)
$K_{0}(\mu)$ $=K^{(0)}_{0}+n^{(0)}_{0}\mu\,$, (174)
$\bar{\kappa}(\mu)$ $=\bar{\kappa}^{(0)}+n^{(0)}_{\bar{\kappa}}\mu\,$, (175)
We want to find a family of solutions that are continuously connected to the $\omega=0$ biconcave disc solution. For this to be the case near $\omega=0$, the solution must have an asymptotic behavior near $\rho=0$ that is compatible with the biconcave disc solution, namely, we will ask for the numerical solution to be compatible with the asymptotic behavior
$\phi(\rho)=-K^{(0)}_{0}\rho\left(c_{1}+\ln\rho\right)+\mathcal{O}\left(\rho^{2}\right)\,$, (176)
for some constant $c_{1}$. For $\phi(\rho)$ to be compatible with (176) the following constraints should hold
$\gamma^{(0)}$ $=\frac{\omega^{2}}{2}\left(n^{(0)}_{\kappa}+n^{(0)}_{\bar{\kappa}}\right)\,$, (177)
$\Delta p$ $=\frac{\omega^{2}}{2}\left(K^{(0)}_{0}n^{(0)}_{\kappa}-\kappa^{(0)}n^{(0)}_{0}\right)\,$. (178)
Constraints (177) and (178) reduce the number of independent thermodynamic coefficients to $\{K^{(0)}_{0},\kappa^{(0)},n^{(0)}_{\kappa},n^{(0)}_{\bar{\kappa}},n^{(0)}_{0}\}$ with $\{K^{(0)}_{0},\kappa^{(0)}\}$ associated to the $\omega=0$ solution and $\{n^{(0)}_{\kappa},n^{(0)}_{\bar{\kappa}},n^{(0)}_{0}\}$ associated to the solution once vorticity is turned on. The exploration of the solutions in this reduced parameter space becomes simpler, and it is the reason we choose to explore this class of solutions. Even then, the full exploration of the parameter space becomes challenging, and we will restrict ourselves to exploring a subspace of it. In particular, we will focus on the following two cases:
case 1: $\qquad n^{(0)}_{0}=0\,,\qquad n^{(0)}_{\bar{\kappa}}=0\,$, (179)
case 2: $\qquad n^{(0)}_{0}=0\,,\qquad n^{(0)}_{\kappa}=0\,$, (180)
In the first case, all thermodynamic parameters are assumed constant with the exception of the bending modulus. In this case, a nonzero surface tension and a nonzero pressure difference are assumed. In the second case, only the effect of a non-constant Gaussian modulus is studied. In this case, a nonzero surface tension and zero pressure difference are assumed. Both cases have similar qualitative behavior while being physically distinct, as one of them requires the presence of a pressure difference. To aid in the discussion, we will define the dimensionless squared vorticities $\hat{\omega}^{2}_{\kappa/\bar{\kappa}}$ as
$\hat{\omega}^{2}_{\kappa}$ $\equiv\frac{r^{2}n^{(0)}_{\kappa}}{\kappa^{(0)}}\omega^{2}\,$, (181)
$\hat{\omega}^{2}_{\bar{\kappa}}$ $\equiv\frac{r^{2}n^{(0)}_{\bar{\kappa}}}{\kappa^{(0)}}\omega^{2}\,$, (182)
These parameters correspond to the dimensionless vorticities square, where the maximum radius $r$, $\kappa^{(0)}$, and $n^{(0)}_{\kappa/\bar{\kappa}}$ were used for this purpose. Note that despite the notation, $\hat{\omega}_{\kappa/\bar{\kappa}}^{2}$ can take negative values when $n^{(0)}_{\kappa/\bar{\kappa}}/\kappa^{(0)}$ takes negative values. For both cases ($n^{(0)}_{\kappa}=0$ or $n^{(0)}_{\bar{\kappa}}=0$), we find two distinct families of solutions that we classify using the dimensionless parameter $rK^{(0)}_{0}$. It is also convenient to use the height $z_{0}=-\int_{0}^{R}d\rho\frac{\phi(\rho)}{\sqrt{1-\phi(\rho)^{2}}}$ to define the aspect ratio $\mathcal{A}$ as
$\mathcal{A}=\frac{z_{0}}{r}\,.$ (183)
Using this definition for $\mathcal{A}$, we will call $\mathcal{A}_{0}$ the aspect ratio at $\hat{\omega}_{\kappa/\bar{\kappa}}^{2}=0$ for any particular given $rK^{(0)}_{0}$. We will use $\mathcal{A}$ to characterize the qualitative features of the solutions. Using $rK^{(0)}_{0}$ the following two distinct families are found:
  1. (a)
    $0\leq rK^{(0)}_{0}<1.2$
    The lower limit $rK^{(0)}_{0}=0$ at $\hat{\omega}_{\kappa/\bar{\kappa}}^{2}=0$ corresponds to the limit in which the biconcave disc solution becomes a sphere; consequently $\mathcal{A}_{0}=1$ for this case. The upper boundary $rK^{(0)}_{0}=1.2$ corresponds to a biconcave disc solution with an aspect ratio of $\mathcal{A}_{0}\approx 0.43$. In figures 1 and 2 a plot for $\mathcal{A}$ vs $\hat{\omega}_{\kappa}^{2}$ and $\mathcal{A}$ vs $\hat{\omega}_{\bar{\kappa}}^{2}$ are shown respectively. These plots are very similar; therefore, figure 2 is presented in Appendix B.
    For all $rK^{(0)}_{0}$ that we looked at and for $\hat{\omega}_{\kappa/\bar{\kappa}}^{2}<0$, the aspect ratio decreases compared to the $\hat{\omega}_{\kappa/\bar{\kappa}}^{2}=0$ solution, namely $\mathcal{A}<\mathcal{A}_{0}$. The shape of the solutions goes from a sphere to a shape that resembles a biconcave disc with aspect ratio $\mathcal{A}$. In between these shapes, and for most of this regime, the membrane takes a shape that resembles an oblate spheroid, see figure (1) for a characteristic example of the shapes found in this regime. It is important to note that while the solution has the visual features of an oblate spheroid, it does not follow the exact analytic form of one.
    These solution follow a decreasing trend in $\mathcal{A}$ as $\hat{\omega}_{\kappa/\bar{\kappa}}^{2}$ decreases until some lower critical value $(\hat{\omega}^{2}_{\kappa/\bar{\kappa}})_{\text{b-crit}}$ that depends on $rK^{(0)}_{0}$, after which we were unable to find solutions that are continuously connected to the trend line connecting the solutions for $(\hat{\omega}^{2}_{\kappa/\bar{\kappa}})_{\text{b-crit}}<\hat{\omega}_{\kappa/\bar{\kappa}}^{2}\leq 0$. Note that $(\hat{\omega}^{2}_{\kappa/\bar{\kappa}})_{\text{b-crit}}$ and its corresponding critical aspect ratio $\mathcal{A}_{\text{b-crit}}$ are different for each $rK^{(0)}_{0}$, although we can note that $\mathcal{A}_{\text{b-crit}}$ is at most within $3\%$ of the aspect ratio for $rK^{(0)}_{0}=1.2$ at $\hat{\omega}_{\kappa/\bar{\kappa}}^{2}=0$.
    For all $rK^{(0)}_{0}$ that we looked at and for $\hat{\omega}_{\kappa/{\bar{\kappa}}}^{2}>0$, the aspect ratio increases compared to the solution $\hat{\omega}_{\kappa/\bar{\kappa}}^{2}=0$, namely $\mathcal{A}>\mathcal{A}_{0}$. The shape of the solutions becomes similar to that of a prolate spheroid, see figure 1 for a characteristic example of the shapes found in this regime. It is important to note that while the solution resembles a prolate spheroid, it does not follow the exact analytic form of one. These solutions follow an increasing trend in $\mathcal{A}$ as $\hat{\omega}_{\kappa/\bar{\kappa}}^{2}$ increases until a critical $(\hat{\omega}^{2}_{\kappa/\bar{\kappa}})_{\text{u-crit}}$ after which we were unable to find solutions that are continuously connected to the trend line connecting the solutions for $0\leq\hat{\omega}_{\kappa/\bar{\kappa}}^{2}<(\hat{\omega}^{2}_{\kappa/\bar{\kappa}})_{\text{u-crit}}$.
  2. (b)
    $1.2\leq rK^{(0)}_{0}\lessapprox 2.42957$
    The lower limit $rK^{(0)}_{0}=1.2$ at $\hat{\omega}_{\kappa/\bar{\kappa}}^{2}=0$ corresponds to a biconcave disc solution with an aspect ratio of $\mathcal{A}_{0}\approx 0.43$. The upper limit $rK^{(0)}_{0}\approx 2.42957$ corresponds to a biconcave disc solution with approximately zero height, namely $\mathcal{A}_{0}\approx 0$. In figure 1 and 2 a plot for $\mathcal{A}$ vs $\hat{\omega}_{\kappa}^{2}$ and $\mathcal{A}$ vs $\hat{\omega}_{\bar{\kappa}}^{2}$ is shown respectively. Note that the plot contains both families, the one currently being discussed and the previously discussed one.
    For all $rK^{(0)}_{0}$ that we looked at and for $\hat{\omega}_{\kappa/\bar{\kappa}}^{2}<0$, the aspect ratio increases compared to the $\hat{\omega}_{\kappa/\hat{\kappa}}^{2}=0$ solution, namely $\mathcal{A}>\mathcal{A}_{0}$. These solutions follow an increasing trend in $\mathcal{A}$ as $\hat{\omega}_{\kappa/\bar{\kappa}}^{2}$ decreases until a critical $(\hat{\omega}^{2}_{\kappa/\bar{\kappa}})_{\text{b-crit}}$ after which we were unable to find solutions that are continuously connected to the trend line connecting the solutions for $(\hat{\omega}^{2}_{\kappa/\bar{\kappa}})_{\text{b-crit}}\leq\hat{\omega}_{\kappa/\bar{\kappa}}^{2}\leq 0$. Note that $(\hat{\omega}^{2}_{\kappa/\bar{\kappa}})_{\text{b-crit}}$ and its corresponding critical aspect ratio $\mathcal{A}_{b-crit}$ are different for each $rK^{(0)}_{0}$, although we can note that $\mathcal{A}_{b-crit}$ is at most within 4% of the aspect ratio for $rK^{(0)}_{0}=1.2$ at $\hat{\omega}_{\kappa/\bar{\kappa}}^{2}=0$. The shape of the solutions remains visually that of a biconcave disc, see figure 1 for a characteristic example of the shapes found in this regime. Note that although the shape resembles a biconcave disc, it is not an exact analytical solution such as (159).
    For all $rK^{(0)}_{0}$ that we looked at and for $\hat{\omega}_{\kappa/\bar{\kappa}}^{2}>0$, the aspect ratio decreases compared to the $\hat{\omega}_{\kappa/\bar{\kappa}}^{2}=0$ solution, namely $\mathcal{A}<\mathcal{A}_{0}$. These solutions follow a decreasing trend in $\mathcal{A}$ as $\hat{\omega}_{\kappa/\bar{\kappa}}^{2}$ increases until a critical $(\hat{\omega}^{2}_{\kappa/\bar{\kappa}})_{\text{u-crit}}$ at which the aspect ratio becomes zero. This suggests a critical rotational speed at which the membrane might suffer a topological change and become a torus, similar to what is observed in active membranes due to an increasing activity [18]. Note that the shapes of the membranes in this regime are still visually that of a biconcave disc.
Notice that the choice of axis in plots 1 and 2, allowed us to present the different types of approximate solutions to membrane shapes in terms of the aspect ratio $\mathcal{A}$. Points in these plots, even the ones with the same symbol (the same value of $rK_{0}^{(0)}$) are not necessarily connected, in the sense that two solutions on the same $rK_{0}^{(0)}$ line do not represent different shapes of the same membrane subject to a different rotation given by $\omega$. For example, if we were interested in the response to rotation of a closed membrane with some given material properties and that is constrained to keep its volume constant, we would need to compute and present the solutions in a different manner.
[…]